OpenStudy (anonymous):

A car dealership has 24 sedans and 18 sports coupes on its lot. What is the ratio of sedans to sports coupes on the lot? A. 3 : 4 B. 2 : 3 C. 3 : 2 D. 4 : 3

OpenStudy (anonymous):

sleepyjess are you there

OpenStudy (sleepyjess):

sedans:sports coupes=24:18 now simplify

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4:3

OpenStudy (sleepyjess):

\(\frak\color{lime}{correct}\)

OpenStudy (anonymous):

i need help with more

OpenStudy (sleepyjess):

Ok, open a new question to ask them.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can i just post it here

OpenStudy (anonymous):

A recipe for waffles calls for 2 cups of flour and 3 teaspoons of baking powder. If Jason uses 4 cups of flour, how much baking powder should he use?

OpenStudy (anonymous):

my answer 6 teaspoons

OpenStudy (sleepyjess):

Yes

OpenStudy (anonymous):

1 more is that ok

OpenStudy (sleepyjess):

yes that one is good

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Which fraction is equivalent to 36%?

OpenStudy (sleepyjess):

what are the fractions?

OpenStudy (anonymous):

a 3 6/10 b 36/10 c 9/25 d 360/100

OpenStudy (anonymous):

my answer 9/25

OpenStudy (sleepyjess):

so the choices are \(3\dfrac{6}{10}~\dfrac{36}{10}~\dfrac{9}{25}~\dfrac{360}{100}\)

OpenStudy (sleepyjess):

yes 9/25 would be correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you for checking my work sleepyjess

OpenStudy (sleepyjess):

No problem!

OpenStudy (anonymous):

bye